determine how many 4-digit numbers can be formed that contains at least one 3.
Answer by sudhanshu_kmr(1152)


total no. of possible 4-digit numbers = 9* 10 * 10 * 10 = 9000

no. of possible 4-digit numbers without using digit 3 = 8*9*9*9 = 5832

so, no. of possible 4-digit numbers containing at least one 3 = 9000-5832
= 3168